OpenStudy (arianna1453):

FAN + MEDAL A paper clip is dropped from the top of a 144‐ft tower, with an initial velocity of 16 ft/sec. Its position function is s(t) = −16t2 + 144. What is its velocity in ft/sec when it hits the ground?

OpenStudy (arianna1453):

@Astrophysics @TrojanPoem @ganeshie8 ??

OpenStudy (astrophysics):

Note the formula is \[s(t) = - g t^2+v_0t+h_0\]

OpenStudy (astrophysics):

So, can you tell me what all the variables mean?

OpenStudy (arianna1453):

No really. Finding velocity confuses me.

ganeshie8 (ganeshie8):

s(t) = −16t2 + 144 this is a position function, i think this represents the height above the ground..

ganeshie8 (ganeshie8):

the height above the ground is 0 when the clip hits the ground, so shouldn't we be solving −16t^2 + 144 = 0 ?

OpenStudy (trojanpoem):

I think you're correct @ganeshie8 , this is the position so , you should solve for 0.

OpenStudy (trojanpoem):

-16 t^2 = -144 t = 3 s v^2 = v0^2 + 2 as v = v0 + at = 16 + 32.174 * 3 = 112.52 ft/s

OpenStudy (arianna1453):

Thats what i thoug. Just set it equal to zero. But the only answer choices are -96 -64 -32 0

OpenStudy (trojanpoem):

@AlexandervonHumboldt2 , Solve this !